OpenStudy (anonymous):

Solve. (3/5)^x=7^1-x

OpenStudy (anonymous):

(3/5)^x = 7^(1 - x) x log (3/5) = (1 - x ) log 7 x log (3/5) = log 7 - x log 7 x [ log (3/5) + log 7 ] = log 7 x [ log (21/5) ] = log 7 x = log 7 / log [ (21/5) ] x = 1.356

OpenStudy (anonymous):

omg thank you!

OpenStudy (anonymous):

No problem. Just here to help!